OpenStudy (anonymous):

The base of a triangle is 4cm greater than the height. The area is 48 cm^2. Find the height and lenght of the base.

jimthompson5910 (jim_thompson5910):

A = (bh)/2 A = ((h+4)h)/2 A = (h(h+4))/2 48 = (h(h+4))/2 48*2 = h(h+4) 96 = h(h+4) 96 = h^2+4h 0 = h^2+4h-96 h^2+4h-96 = 0 (h-8)(h+12)=0 h-8 = 0 or h+12 = 0 h = 8 or h = -12 (throw out negative answer) So the height is 8 cm The base is b = h+4 = 8+4 = 12 cm

OpenStudy (anonymous):

okay thanks !